OpenStudy (anonymous):

solve -2(3-m)+15=6m-4(m-20)

OpenStudy (anonymous):

-2(3-m) = -6 + 2m

OpenStudy (anonymous):

-4(m-20) = -4m + 80

OpenStudy (anonymous):

use both and substitute: -6+2m +15 = 6m-4m+80

OpenStudy (anonymous):

hmm.. Now we end up with 2m on both sides.... this equation is erroneous

OpenStudy (anonymous):

yea that is where i got confused

OpenStudy (anonymous):

there are no solutions

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

ok thanks i thought so just needed a little reassurance

OpenStudy (anonymous):

yw